Pio and Sons Inc. Spends P40, 000 to prepare a bid on a construction project. If the contract is awarded, the estimated revenue will be P300, 000. If the contract is not awarded, the company has a penalty of P50, 000. There is 40% chance that the company will be awarded the contract. Find the EV.


1
Expert's answer
2022-03-20T06:39:50-0400

In case of the contract is awarded, the profit will be P300,000 - P40,000 = P260,000

In case of the contract is not awarded, the losses will be P40,000 + P50,000= P90,000

So EV = 260,000*0.4 - 90,000*(1-0.4) = 50,000